# Learning to utilize image second-order derivative information for crisp edge detection

### v3

@@ IV-B Evaluation methods
To evaluate the quality of generated edge maps, we report the F-score $\left(\frac{2\times Precision\times Recall}{Precision+Recall}\right)$, which is widely used in edge detection [9, 10, 33]. Specifically, $Precision=\frac{TP}{TP+FP}$ and $Recall=\frac{TP}{TP+FN}$, where $TP$, $FP$, and $FN$ represent the number of correctly classified edge pixels, the number of incorrectly classified edge pixels, and the number of missed edge pixels, respectively. Once the network has generated edge maps, the primary step is to apply a threshold to convert them into binary edge maps, and then match these to the groundtruth images to calculate the F-score. There are two choices to set this threshold: optimal dataset scale (ODS) and optimal image scale (OIS). ODS F-score is calculated across the whole dataset by selecting a fixed threshold that maximizes the F-score for all images collectively. OIS F-score is calculated for each image independently, choosing the best threshold for that particular image to maximize its F-score. In addition, we adopt the same strategy as LPCB [33] used to evaluate our method: standard evaluation (S-Eval) and crispness evaluation (C-Eval).
@@
Standard evaluation method. In previous studies, the predicted edge maps first underwent standard post-processing (NMS and morphological thinning), then matched these post-processed edge maps with their groundtruth images to calculate the ODS and OIS. This standard evaluation method is utilized to determine the correctness of predicted edge maps.
@@
Crispness evaluation method. Our purpose is to make the network generate clearer and crisper edge maps. Therefore, we follow the settings in LPCB [33] to evaluate the crispness of our results, which calculate the ODS and OIS without such standard post-processing.
IV-C Ablation experiments on BSDS500 dataset
@@
For the purpose of evaluating each sub-component performance in our model, we conduct a series of ablation experiments on the BSDS500 dataset and provide the analysis in detail. We report S-Eval and C-Eval for the performance evaluation and all these ablation experiments are conducted using MobileNetV2 as the encoder.
Loss function: Firstly, we perform ablation experiments on our hybrid loss function to demonstrate its effectiveness. To identify the optimal fusion of focal loss and focal Tversky loss, we conduct a series of comparison experiments about different $\lambda$, and the comparison results are shown in Table I. It can be observed that, as $\lambda$ decreases, there is an increase in C-Eval; however, S-Eval does not exhibit a corresponding increase. This is because the influence of focal loss decreases. Therefore, we determine the optimal value of $\lambda$ is 0.001. Additionally, when fixed $\lambda$, the best results of both S-Eval and C-Eval are obtained at $\beta=0.7$, thereby the rest ablation experiments are adopted in this setting.
TABLE I: Comparison experiment results for different values of the hyperparameters $\beta$ and $\lambda$ in hybrid focal loss.
hyperparameters S-Eval C-Eval
$\beta$ $\lambda$ ODS OIS ODS OIS
0.7 1 0.805 0.825 0.686 0.694
0.7 0.1 0.807 0.827 0.685 0.694
0.7 0.01 0.809 0.828 0.691 0.696
0.7 0.001 0.805 0.827 0.698 0.705
0.6 0.001 0.802 0.822 0.685 0.691
0.8 0.001 0.803 0.824 0.689 0.694
Network components: The second ablation experiment aims to assess the effectiveness of each proposed component within our model. We first compare the weighted cross-entropy loss with our hybrid loss by training our model using each of these two loss functions separately. As shown in Table III, our hybrid loss function can obtain a higher F-score than the weighted cross-entropy loss in the C-Eval evaluation, which demonstrates the effectiveness of the proposed loss function. When removing individual modules, it is observed that there are varying degrees of decline in both S-Eval and C-Eval. It should be noted that the performance significantly decreases when BRM is excluded, further demonstrating the remarkable power of BRM. In addition to this, we substitute the Laplacian operator in SDMC with the first-order derivative-based Sobel and Scharr operators. The comparative results, presented in Table II, exhibit varying degrees of performance degradation respectively under these alternative setups. This confirms our view that second-order derivative information is better than first-order derivative information. On the other hand, our weighted cross-entropy-based model has a better performance than other models with the same loss function, such as HED and RCF, while having a lower computational cost. These experiment results fully demonstrate the validity of each component.
TABLE II: A comparative analysis of different derivative information.
Methods S-Eval C-Eval
ODS OIS ODS OIS
Sobel 0.803 0.824 0.690 0.697
Scharr 0.803 0.822 0.687 0.693
Laplacian 0.805 0.827 0.698 0.705
TABLE III: Ablation experiment results of each network component. WCE indicates the weighted cross-entropy loss function. HFL indicates the hybrid focal loss function. SDMCM indicates the second-order derivative-based multi-scale contextual enhancement module. BRM indicates the boundary refinement module. DC indicates the dense connection.
Methods WCE HFL SDMCM BRM DC S-Eval C-Eval
ODS OIS ODS OIS
Ours-MobileNetV2 0.812 0.833 0.693 0.702
0.805 0.827 0.698 0.705
0.803 0.822 0.686 0.692
0.794 0.813 0.647 0.654
0.804 0.825 0.687 0.693
@@
Compression ratio analysis: We explore the effect of different channel compression ratios in SDMCM on the model performance and the experiment results are shown in Fig. 6. As the ratio increases, the number of model parameters also increases, resulting in a higher computational cost. However, the performance does not improve with an increasing compression ratio. The best performance is obtained at $r=\frac{1}{4}$. The reason for this phenomenon is that a higher compression ratio causes overfitting, resulting in a lower performance both in S-Eval and C-Eval. Therefore, we set the compression ratio to $\frac{1}{4}$, which achieves a balance between performance and parameters.
Figure 6: The effect of different compression ratios on network performance. From top to down: (a) is S-Eval and (b) is C-Eval. The unit of measurement for ”Parameters” is megabytes (M).
IV-D Comparison to some other SOTA methods
@@
BSDS500: Firstly, we compare our method with some top-performing algorithms on BSDS500. We select a few recent state-of-the-art edge detectors which can be divided into two categories: the first category is methods without deep learning, which includes Canny [27], gPb-UCM [12] and SE [29]; the second category is approaches using deep learning technique, which including DeepContour [47], DeepEdge [48], HED [9], RCF [10], BDCN [14], CED [32], LPCB [33], DRC [15], PiDiNet [18], EDTER [41] and CHRNet [11]. We additionally reference the studies of other researchers who utilize extra training data sourced from the PASCAL VOC Context dataset [49] and adopt multi-scale testing. Table IV presents the ODS and OIS F-score in S-Eval and C-Eval. Fig. 7 shows some qualitative comparison results, and Fig. 8 draws the Precision-Recall curves.
Figure 7: Some examples from state-of-the-art (SOTA) methods on BSDS500. From left to right: (a) is the original image, (b) is the groundtruth picture, (c), (d), (e), and (f) are the predictions of HED, RCF, CED, and DRC, respectively. (g) is the output edge map from ours-EfficientV2.
@@
As shown in Fig. 7, HED and RCF are two classical algorithms for edge detection using VGG-16 [43] architecture, while their edge maps are noisy and blurred. CED and DRC share the same goal as ours, which aims to generate crisp edge maps without any post-processing. However, compared to the other two methods, our method can generate higher-quality and crisper edge maps. Specifically, in the third row of Fig. 7, the contours of the man’s eyes and nose are noisy, but our method can filter out these pixels, unlike the DRC which incorrectly classifies them. Similar results can be observed in other examples.
@@
TABLE IV: Quantitative evaluation of comparison on BSDS500 dataset. VOC indicates training with extra PASCAL VOC Context data. MS indicates the multi-scale testing.
Methods S-Eval C-Eval
ODS OIS ODS OIS
Canny [27] 0.611 0.676 - -
gPb-UCM [12] 0.729 0.755 - -
SE [29] 0.743 0.764 - -
DeepEdge [48] 0.753 0.772 - -
DeepContour [47] 0.757 0.776 - -
HED [9] 0.788 0.808 0.576 0.591
DRC [15] 0.802 0.818 0.697 0.705
CED [32] 0.794 0.811 0.642 0.656
RCF [10] 0.798 0.815 0.585 0.604
LPCB [33] 0.800 0.816 0.693 0.700
BDCN [14] 0.806 0.826 0.636 0.650
PiDiNet [18] 0.789 0.803 0.578 0.587
EDTER [41] 0.824 0.841 0.698 0.706
CHRNet [11] 0.787 0.788 - -
Ours-ShuffleNetV2 0.790 0.812 0.670 0.677
Ours-MobileNetV2 0.805 0.827 0.698 0.705
Ours-EfficientNetV2 0.826 0.846 0.720 0.726
Ours-ShuffleNetV2-VOC 0.804 0.825 0.683 0.691
Ours-MobileNetV2-VOC 0.815 0.835 0.705 0.711
Ours-EfficientNetV2-VOC 0.827 0.846 0.717 0.723
Ours-ShuffleNetV2-VOC-MS 0.813 0.833 0.672 0.677
Ours-MobileNetV2-VOC-MS 0.821 0.842 0.684 0.690
Ours-EfficientNetV2-VOC-MS 0.829 0.851 0.694 0.700
@@
Figure 8: Precision-Recall curves on BSDS500 dataset. Our method achieves the best performance (ODS=0.829).
@@
Table IV clearly shows our method using ShuffleNetV2 still achieves a competitive performance when trained with only the BSDS500 dataset, which outperforms HED and RCF in both S-Eval and C-Eval. When using the MobileNetV2 as the encoder, there is an increment in S-Eval and C-Eval whose ODS F-scores are improved from 0.790 to 0.805 and 0.670 to 0.698, respectively. The highest performance is achieved by using the EfficientNetV2 with the ODS=0.826 and OIS=0.846 in S-Eval. These results are better than those of EDTER, which is the first edge detection algorithm using Transformer, particularly considering our model has fewer parameters. In the standard evaluation, when we mix the extra PASCAL VOC Context data into the BSDS500 dataset, there is a significant improvement in the ODS and OIS. The corresponding values are 0.829 and 0.851, reaching the SOTA performance among all the methods. In the crispness evaluation, compared to the CED, LPCB, and DRC, our method using EfficientNetV2 trained with only the BSDS500 dataset demonstrates a significant performance margin, the ODS and OIS values are 0.720 and 0.726, respectively. Specifically, our ODS and OIS are 12.1% and 10.7% higher than those of CED, 3.9% and 3.7% higher than those of LPCB, 3.3% and 3.0% higher than those of DRC. These quantitative results are consistent with the qualitative visualization. All the comparison results demonstrate that our method has excellent performance, we successfully address the issue of edge thickness while improving the method’s performance. Fig. 8 shows that the performance of the human eye in edge detection is 0.803. Our results are better than the human level and obtain the best performance among the current SOTA methods.
@@
NYUD-V2: Secondly, we select the NYUD-V2 dataset to conduct another set of comparison experiments. We adopt some methods as before, which consists of algorithms without using deep learning such as gPb-UCM [12], OEF [50], gPb+NG [51], SE [29] and SE+NG+ [52], and recent top edge detectors based on deep learning such as HED [9], RCF [10], BDCN [14], DRC [15], and CHRNet [11].
Figure 9: Precision-Recall curves on NYUD-V2 dataset. Our method trained on the RGB data and the HHA data achieves a top performance (ODS=0.768).
Figure 10: Some examples from state-of-the-art (SOTA) methods on NYUD-V2. From left to right: (a) is the RGB image, (b) is the HHA image, (c) is the groundtruth picture, (d) is the prediction of DRC, and (e) is the generated edge map of our method. Both the predictions of DRC and our method are only on the RGB image.
@@
Since NYUD-V2 dataset consists of two types of images: RGB images and HHA images. We train and test our model on three versions following the previous researchers’ works: (a) RGB images only (Ours-RGB); (b) HHA images only (Ours-HHA); (c) directly averaging the predictions from the RGB version and the HHA version (Ours-RGB-HHA). Some predicted examples are shown in Fig. 10. The quantitative evaluation results are summarized in Table V and the Precision-Recall curves are drawn in Fig. 9.
TABLE V: Quantitative evaluation values of comparison on NYUD-V2 dataset. RGB indicates RGB images. HHA indicates HHA images. RGB-HHA indicates averaging the prediction of RGB images and HHA images. Ours indicates our method based on EfficientNetV2.
Methods S-Eval C-Eval
ODS OIS ODS OIS
gPb-UCM [12] 0.631 0.661 - -
OEF [50] 0.651 0.667 - -
gPb+NG [51] 0.687 0.716 - -
SE [29] 0.695 0.708 - -
SE+NG+ [52] 0.706 0.734 - -
HED-RGB [9] 0.722 0.737 0.387 0.404
HED-HHA [9] 0.691 0.704 0.335 0.350
HED-RGB-HHA [9] 0.746 0.764 0.368 0.384
RCF-RGB [10] 0.745 0.749 0.395 0.412
RCF-HHA [10] 0.701 0.702 0.333 0.348
RCF-RGB-HHA [10] 0.764 0.778 0.374 0.397
BDCN-RGB [14] 0.728 0.762 0.414 0.439
BDCN-HHA [14] 0.704 0.716 0.347 0.367
BDCN-RGB-HHA [14] 0.766 0.779 0.375 0.392
DRC-RGB [15] 0.749 0.762 0.411 0.455
DRC-RGB-HHA [15] 0.711 0.722 0.370 0.382
DRC-RGB-HHA [15] 0.769 0.782 0.403 0.436
CHRNet-RGB [11] 0.729 0.745 - -
CHRNet-HHA [11] 0.718 0.731 - -
CHRNet-RGB-HHA [11] 0.750 0.774 - -
Ours-RGB 0.757 0.768 0.546 0.559
Ours-HHA 0.717 0.727 0.435 0.450
Ours-RGB-HHA 0.768 0.780 0.513 0.524

### v4

@@ 4.2 Evaluation methods
To evaluate the quality of generated edge maps, we report the F-score $\left(\frac{2\times Precision\times Recall}{Precision+Recall}\right)$ and Average Precision (AP), which are widely used in edge detection [24, 25, 48]. Specifically, $Precision=\frac{TP}{TP+FP}$ and $Recall=\frac{TP}{TP+FN}$, where $TP$, $FP$, and $FN$ represent the number of correctly classified edge pixels, the number of incorrectly classified edge pixels, and the number of missed edge pixels, respectively. Once the network has generated edge maps, the primary step is to apply a threshold to convert them into binary edge maps, and then match these to the groundtruth images to calculate the F-score. There are two choices to set this threshold: optimal dataset scale (ODS) and optimal image scale (OIS). ODS F-score is calculated across the whole dataset by selecting a fixed threshold that maximizes the F-score for all images collectively. OIS F-score is calculated for each image independently, choosing the best threshold for that particular image to maximize its F-score. AP summarizes the Precision-Recall curve by calculating the area under this curve, thereby providing a single score that captures the model’s performance across all possible thresholds. In addition, we also report the number of parameters (Params) and Frames Per Second (FPS) to evaluate the efficiency of the model.
To comprehensively evaluate the crispness of edge maps, following these previous works [47, 48], we adopt the same strategy to evaluate our method including standard evaluation (S-Eval) and crispness evaluation (C-Eval).
@@
Standard evaluation method. This method first applies standard post-processing process (NMS and morphological thinning) to the predicted edge maps, then matches these post-processed results with groundtruth images to calculate ODS, OIS, and AP. The S-Eval aims to assess the correctness of predicted edge maps.
@@
Crispness evaluation method. To evaluate whether the network can directly generate clean and crisp edge maps, we calculate ODS, OIS, and AP on the raw predictions without such standard post-processing steps.
4.3 Ablation experiments on BSDS500 dataset
@@
To analyze the effectiveness of each configuration in our LUS-Net, we conduct a series of ablation experiments on the BSDS500 dataset with MobileNetV2 as the encoder. The performance is evaluated using both S-Eval and C-Eval methods.
Loss function: Firstly, we perform ablation experiments on various loss functions to demonstrate the effectiveness of our HFL. To identify the optimal configuration of HFL, we conduct a series of comparison experiments with different parameters, and the comparison results are shown in rows 1-6 of Table 1. It can be observed that, as $\lambda$ decreases from 1 to 0.001, there is a notable increase in C-Eval metrics, while S-Eval metrics maintain relatively stable performance. Such observation can be explained by the complementary roles of the two loss components: focal Tversky loss serves to regulate the crispness of edge delineation, whereas focal loss ensures the accuracy of pixel-level classification in the predicted edge maps. Therefore, the optimal value of $\lambda$ is 0.001. Additionally, when fixed $\lambda=0.001$, the best results of both S-Eval and C-Eval are obtained at $\beta=0.7$, outperforming other values (0.6 and 0.8). In rows 11-14, it can be observed that the best performance is obtained at $\gamma=0.75$. Furthermore, we compare our HFL loss with other loss functions, including Weighted Cross-Entropy [24] (WCE), Hybrid Quality Focal Loss [65] (HQFL), and Hybrid Varifocal Loss [66] (HVFL). The comparison results shown in rows 7-10 demonstrate that the HFL maintains competitive performance in S-Eval (ODS=0.805, OIS=0.827, AP=0.805). More importantly, HFL demonstrates superior performance across all metrics in C-Eval (ODS=0.698, OIS=0.705, AP=0.754), notably outperforming WCE, HQFL, and HVFL. These results strongly indicate that our proposed HFL effectively enhances the model’s capability to produce crisp and clean edge maps while maintaining robust performance in pixel-level classification.
Table 1: Ablation experiment results about different configurations in our network.
Method Compression ratio Loss Derivative S-Eval C-Eval
LUS-Net-M ODS OIS AP ODS OIS AP
0.25 HFL$\left(\beta=0.7,\lambda=1\right)$ Laplacian 0.805 0.825 0.807 0.686 0.694 0.742
0.25 HFL$\left(\beta=0.7,\lambda=0.1\right)$ Laplacian 0.807 0.827 0.807 0.685 0.694 0.744
0.25 HFL$\left(\beta=0.7,\lambda=0.01\right)$ Laplacian 0.809 0.828 0.808 0.691 0.696 0.750
0.25 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.805 0.827 0.805 0.698 0.705 0.754
0.25 HFL$\left(\beta=0.6,\lambda=0.001\right)$ Laplacian 0.802 0.822 0.803 0.685 0.691 0.740
0.25 HFL$\left(\beta=0.8,\lambda=0.001\right)$ Laplacian 0.803 0.824 0.804 0.689 0.694 0.740
0.25 HFL$\left(\gamma=0.5,\lambda=0.001\right)$ Laplacian 0.807 0.825 0.814 0.682 0.690 0.742
0.25 HFL$\left(\gamma=0.75,\lambda=0.001\right)$ Laplacian 0.805 0.827 0.805 0.698 0.705 0.754
0.25 HFL$\left(\gamma=1.0,\lambda=0.001\right)$ Laplacian 0.806 0.826 0.814 0.675 0.691 0.728
0.25 HFL$\left(\gamma=1.25,\lambda=0.001\right)$ Laplacian 0.805 0.827 0.821 0.682 0.689 0.735
0.25 WCE Laplacian 0.812 0.833 0.850 0.693 0.702 0.745
0.25 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.805 0.827 0.805 0.698 0.705 0.754
0.25 HQFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.803 0.821 0.808 0.682 0.694 0.727
0.25 HVFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.805 0.825 0.804 0.690 0.698 0.753
0.125 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.801 0.821 0.810 0.682 0.687 0.739
0.25 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.805 0.827 0.805 0.698 0.705 0.754
0.5 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.802 0.824 0.798 0.692 0.698 0.741
0.25 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Laplacian 0.805 0.827 0.805 0.698 0.705 0.754
0.25 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Sobel 0.803 0.824 0.805 0.690 0.697 0.745
0.25 HFL$\left(\beta=0.7,\lambda=0.001\right)$ Scharr 0.803 0.822 0.803 0.687 0.693 0.742
@@
Compression ratio: We also explore the effect of different channel compression ratios $r$ in SDMCM on the model performance and the experiment results are shown in rows 11-13 of Table 1 and Fig. 7. As the ratio increases, the number of model parameters also increases, resulting in a higher computational cost. However, the performance does not improve with an increasing compression ratio. The best performance is obtained at $r=\frac{1}{4}$. The reason for this phenomenon is that a higher compression ratio causes overfitting, resulting in a lower performance both in S-Eval and C-Eval. Therefore, we set the compression ratio to $\frac{1}{4}$, which achieves a balance between performance and parameters.
Figure 7: The effect of different compression ratios on network performance.
Image derivative information: In addition, we investigate the effectiveness of different derivative operators by replacing the Laplacian operator in SDMCM with first-order derivative operators (Sobel and Scharr). As shown in rows 14-16 of Table 1, both alternative operators lead to consistent performance degradation across all metrics, substantiating our theoretical analysis that second-order derivative information provides more discriminative features for edge detection compared to first-order derivatives.
Network components: Finally, we conduct the last ablation studies to evaluate the contribution of each proposed component within our network, and the results are presented in Table 2. The experiment results demonstrate that our complete model with all components achieves the best performance. When removing each component from our model, the performance shows varying degrees of degradation in both S-Eval and C-Eval. Notably, the exclusion of BRM leads to the most significant performance drop, underscoring its crucial role in our framework. These comprehensive experiments convincingly demonstrate the effectiveness and necessity of each proposed component.
Table 2: Ablation experiment results of each network component. DC indicates the dense connection.
Methods SDMCM BRM DC S-Eval C-Eval
ODS OIS AP ODS OIS AP
LUS-Net-M 0.805 0.827 0.805 0.698 0.705 0.754
0.803 0.822 0.801 0.686 0.692 0.736
0.794 0.813 0.791 0.647 0.654 0.675
0.804 0.825 0.811 0.687 0.693 0.739
4.4 Comparison with SOTA methods
@@
BSDS500: Firstly, we compare our method with top-performing algorithms on BSDS500. We select recent state-of-the-art edge detectors which can be divided into two categories: the first category is methods without deep learning, which includes Canny [42], gPb-UCM [27], SE [44], SCG [67], PMI [68], OEF [69], MES [70], and Liu et al. [71]; the second category is approaches using deep learning technique, which including DeepContour [72], DeepEdge [73], HED [24], RCF [25], BDCN [29], CED [47], LPCB [48], DRC [30], PiDiNet [33], CATS [54], FCL-Net [55], EDTER [57], Boosting-ED [74], CHRNet [26], DexiNed [45], BetaNet [75], GCB-Net [56], EdgeSAM [76], MuGE [59], and RankED [60]. We additionally reference the studies of prior researchers who utilize extra training data sourced from the PASCAL VOC Context dataset [77] and adopt multi-scale testing to further improve the performance. Table 3 presents all metrics including S-Eval, C-Eval, Params, and FPS. Fig. 8 shows some qualitative comparison results, and Fig. 9 draws the Precision-Recall curves in S-Eval.
Figure 8: Some examples from SOTA methods on BSDS500.
@@
As shown in Fig. 8, HED and RCF are two classical algorithms for edge detection using VGG-16 [61] architecture, while their edge maps are noisy and blurred. CED and DRC share the same goal as ours, which aims to generate crisp edge maps without any post-processing. However, compared to the other two methods, our method can generate higher-quality and crisper edge maps. Specifically, in the third row of Fig. 8, the contours of the man’s eyes and nose are false edge pixels because they are not labeled in the corresponding groundtruth image, but our method can filter out these pixels, unlike the DRC which incorrectly classifies them. Similar results can be observed in other examples.
@@
Table 3: Quantitative comparison on BSDS500 dataset. $\dagger$ indicates training with extra PASCAL VOC Context data. $\ddagger$ indicates the multi-scale testing. -S, -M, and -E respectively refers to ShuffleNetV2, MobileNetV2, and EfficientNetV2. $\star$ denotes the cited GPU speed, and * denotes the inference speed tested on A40 GPU.
Methods S-Eval C-Eval Params FPS
ODS OIS AP ODS OIS AP
Canny [42] 0.611 0.676 0.520 - - - - -
gPb-UCM [27] 0.729 0.755 0.745 - - - - -
SE [44] 0.743 0.764 0.800 - - - - -
SCG [67] 0.739 0.758 0.773 - - - - -
PMI [68] 0.741 0.769 0.799 - - - - -
OEF [69] 0.746 0.770 0.820 - - - - -
MES [70] 0.756 0.776 0.756 - - - - -
Liu et al. [71] 0.614 - - - - - - -
DeepEdge [73] 0.753 0.772 0.807 - - - - -
DeepContour [72] 0.757 0.776 0.790 - - - - -
HED [24] 0.788 0.808 0.840 0.576 0.591 - 14.7M 30$\star$
RCF [25] 0.798 0.815 - 0.585 0.604 - 14.8M 30$\star$
DRC [30] 0.802 0.818 0.800 0.697 0.705 - 23.5M 30*
CED [47] 0.794 0.811 - 0.642 0.656 - 14.9M -
[…]
LPCB [48] 0.800 0.816 - 0.693 0.700 - - -
BDCN [29] 0.806 0.826 0.847 0.636 0.650 - 16.3M 44$\star$
PiDiNet [33] 0.789 0.803 - 0.578 0.587 - 0.71M 69*
CATS [54] 0.800 0.816 - 0.666 0.676 - - -
FCL-Net [55] 0.807 0.822 - 0.652 0.666 0.570 16.46M 7.8*
EDTER [57] 0.824 0.841 0.880 0.698 0.706 - 468.84M 2.1$\star$
[…]
CHRNet [26] 0.787 0.788 0.801 - - - - -
DexiNed [45] 0.729 0.745 0.583 - - - 35.2M -
BetaNet [75] 0.803 0.822 - - - - - -
GCB-Net [56] 0.811 0.828 0.809 0.678 0.686 0.731 14.97M 61*
EdgeSAM [76] 0.838 0.852 0.893 - - - - -
MuGE [59] 0.850 0.856 0.896 0.671 0.678 0.713 93.9M 28*
RankED [60] 0.824 0.840 0.895 0.611 0.619 0.653 114.4M -
LUS-Net-S 0.790 0.812 0.800 0.670 0.677 0.640 2.77M 81.17*
[…]
LUS-Net-M 0.805 0.827 0.805 0.698 0.705 0.754 11.51M 76.14*
[…]
LUS-Net-E 0.826 0.846 0.820 0.720 0.726 0.779 70.41M 43.47*
[…]
@@
Figure 9: Precision-Recall curves on BSDS500 and NYUD-V2 dataset, respectively.
@@
Table 3 clearly shows that our LUS-Net-S still achieves a competitive performance when trained with only the BSDS500 dataset, outperforming HED, RCF, and PiDiNet in both S-Eval and C-Eval. When using the MobileNetV2 as the encoder, there are improvements in S-Eval and C-Eval, with ODS F-scores improving from 0.790 to 0.805 and 0.670 to 0.698, respectively. The highest performance is achieved by LUS-Net-E with ODS=0.826, OIS=0.846, and AP=0.820 in S-Eval. Compared to both EDTER, the first Transformer-based edge detection method, and the recent SOTA method RankED, our method demonstrates superior performance while maintaining a more efficient architecture with significantly fewer parameters (70.41M vs. 468.84M for EDTER and 114.4M for RankED). In addition, when we mix the extra PASCAL VOC Context data into the BSDS500 dataset, we observe a significant improvement in the S-Eval metrics (ODS=0.829, OIS=0.851, and AP=0.846), achieving competitive performance among all the methods. In the crispness evaluation, compared to existing methods addressing the edge thickness issue, including CED, LPCB, DRC, and GCB-Net, our LUS-Net-E demonstrates a significant performance margin, with ODS, OIS, and AP values of 0.720, 0.726, and 0.779 respectively, achieving SOTA performance among all the methods. Specifically, our ODS and OIS are 12.1% and 10.7% higher than those of CED, 3.9% and 3.7% higher than those of LPCB, 3.3% and 3.0% higher than those of DRC, and 6.2% and 5.8% higher than those of GCB-Net. These quantitative results are consistent with the qualitative visualization. It is worth noting that although our approach employs a single-granularity architecture in contrast to MuGE’s multi-granularity design, and thus shows marginally lower performance under S-Eval, it exhibits remarkably better performance in C-Eval (0.720 vs. 0.611). In terms of model efficiency, our approach demonstrates impressive performance in both parameter efficiency and inference speed. Notably, LUS-Net-S achieves competitive performance with only 2.77M parameters while maintaining a high inference speed of 81.17 FPS on A40 GPU, surpassing previous crisp edge detection methods such as DRC [30] (23.5M, 30 FPS) and GCB-Net [56] (14.97M, 61 FPS). The larger variant LUS-Net-M (11.51M parameters, 76.14 FPS) and LUS-Net-E (70.41M parameters, 43.47 FPS) also maintain efficient parameter-performance trade-offs. All these comparison results demonstrate that our method has excellent performance, we successfully address the issue of edge thickness while improving the overall performance. Fig. 9 shows that the performance of the human eye in edge detection is 0.803. Our results surpass the human level and achieve the best performance among current SOTA methods.
@@
NYUD-V2: Secondly, we select the NYUD-V2 dataset to conduct another set of comparison experiments. We adopt some methods as before, which consists of algorithms without using deep learning such as gPb-UCM [27], OEF [69], gPb+NG [78], SE [44] and SE+NG+ [79], and recent top edge detectors based on deep learning such as HED [24], RCF [25], BDCN [29], CATS [54], DRC [30], PiDiNet [33], DexiNed [45], LPCB [48], CHRNet [26], BetaNet [75], GCB-Net [56], EDTER [57], Boosting-ED [74], RankED [60], and EdgeSAM [76].
Figure 10: Some examples from SOTA methods on NYUD-V2.
@@
Since NYUD-V2 dataset consists of two types of images: RGB images and HHA images. We train and test our model on three versions following the previous researchers’ works: (a) RGB images only; (b) HHA images only; (c) directly averaging the predictions from the RGB version and the HHA version. Some predicted examples are shown in Fig. 10. The quantitative evaluation results are summarized in Table 4 and the Precision-Recall curves are drawn in Fig. 9.
Table 4: Quantitative comparison results on NYUD-V2 dataset. $\dagger$ refers to training with extra BSDS500 data. $\star$ indicates the cited GPU speed, and * denotes the inference speed tested on A40 GPU.
Methods S-Eval Params FPS
RGB HHA RGB-HHA
ODS OIS AP ODS OIS AP ODS OIS AP
gPb-UCM [27] 0.631 0.661 0.562 - - - - - - -
OEF [69] 0.651 0.667 0.653 - - - - - - -
gPb+NG [78] 0.687 0.716 0.629 - - - - - - - -
SE [44] 0.695 0.708 0.719 - - - - - - - -
SE+NG+ [79] 0.706 0.734 0.549 - - - - - - - -
HED [24] 0.720 0.734 0.734 0.682 0.695 0.702 0.746 0.761 0.786 14.7M 20$\star$
RCF [25] 0.729 0.742 - 0.705 0.715 - 0.757 0.771 - 14.8M 20$\star$
BDCN [29] 0.748 0.763 0.770 0.707 0.719 0.731 0.765 0.781 0.813 16.3M 44$\star$
CATS [54] 0.732 0.746 - 0.693 0.703 - 0.755 0.770 - - -
DRC [30] 0.749 0.762 0.718 0.711 0.722 0.677 0.769 0.782 0.771 23.5M 30*
PiDiNet [33] 0.733 0.747 - 0.715 0.728 - 0.756 0.773 - 0.71M 61*
DexiNed [45] 0.658 0.674 0.556 - - - - - - 35.2M -
LPCB [48] 0.739 0.754 - 0.707 0.719 - 0.762 0.778 - 15.7M 30$\star$
CHRNet [26] 0.729 0.745 - 0.718 0.731 - 0.750 0.774 - - -
BetaNet [75] 0.734 0.748 - 0.714 0.728 - - - -
GCB-Net [56] 0.751 0.765 0.719 0.708 0.720 0.671 0.766 0.782 0.777 14.97M 50*
EDTER [57] 0.774 0.789 0.797 0.703 0.718 0.727 0.780 0.797 0.814 468.84M 2.1$\star$
[…]
RankED [60] 0.780 0.793 0.826 - - - - - - 114.4M -
EdgeSAM [76] 0.783 0.797 0.805 - - - - - - - -
LUS-Net 0.757 0.768 0.705 0.717 0.726 0.644 0.768 0.780 0.775 70.41M 34.26*
Methods C-Eval Params FPS
RGB HHA RGB-HHA
ODS OIS AP ODS OIS AP ODS OIS AP
HED [24] 0.387 0.404 - 0.335 0.350 - 0.368 0.384 - 14.7M 20$\star$
RCF [25] 0.395 0.412 - 0.333 0.348 - 0.374 0.397 - 14.8M 20$\star$
BDCN [29] 0.414 0.439 - 0.347 0.367 - 0.375 0.392 - 16.3M 44$\star$
CATS [54] 0.405 0.418 - 0.349 0.362 - 0.382 0.397 - - -
PiDiNet [33] 0.399 0.424 0.344 - - - - - - 0.71M 61*
DRC [30] 0.403 0.416 - 0.370 0.382 - 0.403 0.436 - 23.5M 30*
GCB-Net [56] 0.456 0.468 0.404 0.391 0.404 0.337 0.420 0.433 0.380 14.97M 50*
EDTER [57] 0.416 0.438 0.372 0.394 0.410 0.348 0.402 0.420 0.368 468.84M 2.1$\star$
RankED [60] 0.477 0.490 0.474 - - - - - - 114.4M -
LUS-Net 0.540 0.554 0.508 0.435 0.450 0.400 0.513 0.524 0.501 70.41M 34.26*